QS=QD

-20P + 2P = 100 - 2P

4P = 120

4P / 120 

P = 30

Q = 40

(a)  Graph the supply and demand curves. Be sure to calculate the P and Q intercepts for demand and the P intercept for supply. Calculate and illustrate the equilibrium price and quantity. [Hint: Show your work. (b)  Calculate both the demand and supply elasticity around the equilibrium point. [Hint: you can use either the point method or the average arc (midpoint) method. (c)  Suppose the government implements a price ceiling of $20/unit in this market. Is this price ceiling binding on the market? What are the quantities demanded and supplied at the price ceiling? How many units are exchanged at this price? 

Given the effects of the policy, is there a potential for illegal trade? Briefly explain your answers where necessary.


1
Expert's answer
2021-10-06T09:44:25-0400

At equilibrium Qs"=" Ad

"-" 20"+" 2P"=100-2P"

4P"=" 120

P"=" 30

Q"=" 40

For the graph when Qs"=" 0

0"=" "-20+2P"

P"=" 10

When Qd"=" 0

0"=" 100"-2P"

P"=" 50





b. Demand and supply elasticities around equilibrium point can be found as derivatives from supply and demand curves. That is Ed"=" "-2," Es"=2" . Therefore demand is inelastic and supply is elastic.

c. If government imposes a price ceiling of $20 per unit then:

Qs"=" Qd

100"-" 2"(P-20)" "=" "-20+2P"

160"=" 4P

P"=" 40

Qd"=" "-20+2(40)=" 60 units

Price ceiling will be non-binding since  the level of the price ceiling is greater than equilibrium price.

Price is above equilibrium price. Quantity supplied will exceed quantity demanded and excess surplus will result.

Governments impose price ceilings in order to keep the price of some necessary goods or services affordable.

A price ceiling is a mega maximum price that one pays for some goods or services.